OpenStudy (anonymous):

identify the initial amount a and the growth factor b in each exponential function. f(t)=1.5^t

OpenStudy (zehanz):

Initial amount is outcome of formula if t=0. Growth factor: factor by which the amount is multiplied every time t is incremented 1, so: b=f(t+1)/f(t), e.g. b=f(2)/f(1). Once you've found b in this way, you'll notice somthing...